One positive number is 1/5 of another number. The difference between the two numbers is 76. Find the numbers.
Found 2 solutions by josgarithmetic, rfer:
Answer by josgarithmetic(39838)

Let p be the unknown positive number.
p-(1/5)p=76
p(1-(1/5))=76
(4/5)p=76
p=5*76/4
p=95

Answer by rfer(16322)

x-1/5x=76
4/5x=76
x=76/(4/5)
x=76*5/4
x=95
-----
95-1/5*95=76
95-19=76